# Goal-Oriented Status Updating for Real-time Remote Inference over Networks with Two-Way Delay

### v3

@@ VI Conclusion
This paper studies a goal-oriented communication design problem for remote inference, where an intelligent neural network model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location. We derive two optimal scheduling policies under time-invariant and time-variable packet length selection. These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory. Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one sixth with respect to age-based scheduling of unit-length packets.
@@ Acknowledgment
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big{(}\delta,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ evolves according to equation (1). Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@ Appendix B Proof of Theorem 1
Step 2: We construct a new SMDP based on the results obtained in the first step. In the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. At each decision time $D_{i+1}$, the system state is represented only by the delay state $c_{i}$ in the $i$-th epoch. We fix the waiting time decisions by using optimal threshold rule $\tau_{l}\big{(}\delta,c_{i}\big{)}$, and the action involves determining the buffer position $b_{i+1}$. A detailed description of this SMDP can be found in Appendix D. The Bellman optimality equation of this new SMDP is given by equation (III-B).
@@
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big{(}\delta,d,c_{i}\big{)}$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ and the length $d(t)$ of the packet most recently delivered by time slot $t$ evolve according to equations (1) and (2), respectively. Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@
(iii) State and State Transitions: The system state at each decision time $D_{i+1}$ is the delay state $C_{i}=c_{i}$ in the $i$-th epoch, which evolves according to the finite-state ergodic Markov chain defined in Section (II). The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.

### v4

@@ VI Conclusion
This paper studies a goal-oriented communication design problem for remote inference, where an intelligent neural network model on the receiver side predicts the real-time value of a target signal using data packets transmitted from a remote location. We derive two optimal scheduling policies under time-invariant and time-variable packet length selection. These policies minimize the expected time-average inference error for a possibly non-monotonic inference error function by considering the interplay between packet length and transmission delay, as well as by exploiting delay memory. Finally, through the simulations, we demonstrate that our goal-oriented scheduler drops inference error down to one-sixth with respect to age-based scheduling of unit-length packets.
@@ Acknowledgment
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big(\delta,c_{i}\big)$, where $\delta=\Delta(A_{i})$ and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ evolves according to equation (1). Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section II. The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@ Appendix B Proof of Theorem 1
Step 2: We construct a new SMDP based on the results obtained in the first step. In the new SMDP, the decision time is $D_{i+1}$ instead of $A_{i}$. At each decision time $D_{i+1}$, the system state is represented only by the delay state $c_{i}$ in the $i$-th epoch. We fix the waiting time decisions by using the optimal threshold rule $\tau_{l}\big(\delta,c_{i}\big)$, and the action involves determining the buffer position $b_{i+1}$. A detailed description of this SMDP can be found in Appendix D. The Bellman optimality equation of this new SMDP is given by equation (III-B).
@@
(iii) State and State Transitions: The system state at each decision time $A_{i}$ is represented by the tuple $\big(\delta,d,c_{i}\big)$, where $\delta=\Delta(A_{i})$, $d=d(A_{i})$, and $c_{i}$ is the delay state in the $i$-th epoch. The AoI $\Delta(t)$ at time slot $t$ and the length $d(t)$ of the packet most recently delivered by time slot $t$ evolve according to equations (1) and (2), respectively. Meanwhile, the delay state $C_{i}=c_{i}$ in the $i$-th epoch evolves according to the finite-state ergodic Markov chain defined in Section II. The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.
@@
(iii) State and State Transitions: The system state at each decision time $D_{i+1}$ is the delay state $C_{i}=c_{i}$ in the $i$-th epoch, which evolves according to the finite-state ergodic Markov chain defined in Section II. The Markov chain makes a single transition at each decision time $A_{i}$ and none otherwise.